OpenStudy (johanna1736):

Help please! Solve for x: |x − 2| + 10 = 12

OpenStudy (johanna1736):

x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution

OpenStudy (johanna1736):

I already solved one and got x=0 as an answer.

OpenStudy (lightlegendx):

x=4

MsBrains (ms-brains):

What are you options?

OpenStudy (lightlegendx):

I can assume those are brackets

OpenStudy (anonymous):

x=4 and x=0

OpenStudy (johanna1736):

I'm not sure if that means no solution or... I posted them above.

OpenStudy (lightlegendx):

Therefore, giving you x=4

MsBrains (ms-brains):

The answer is x = 0 and x = 4

OpenStudy (johanna1736):

Thanks!!!

OpenStudy (lightlegendx):

x = 4 ( 4 - 2 ) + 10 = 12

MsBrains (ms-brains):

Welcome. c;

OpenStudy (anonymous):

that is not brackets it is modulus function

OpenStudy (anonymous):

\[(x-2) + 10 =12 , x \ge 2\] gives you x=4 \[-(x-2)+10=12 , x <2\] gives you x=0